In 6 hours Renu walks 6 km more than Kopila walks in 3 hours and in 8 hours Kopila walks 12 km more than renu walks in 9 hours .Find their speed in km per hour.
Answer by lwsshak3(11628)

In 6 hours Renu walks 6 km more than Kopila walks in 3 hours and in 8 hours Kopila walks 12 km more than renu walks in 9 hours .Find their speed in km per hour.
***
let x=Renu's walking speed
let y=Kopila's walking speed
distance=travel time*speed
..
6x-3y=6
8y-9x=12
..
6x-3y=6
-9x+8y=12
..
48x-24y=48
-27x+24y=36
add
21x=84
x=3
3y=6x-6=18-6=12
Renu's walking speed=3 km/hr
Kopila's walking speed=12 km/hr